OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = 2x2 – 8x + 6? Answer A. x–intercepts: (–3, 0) and (–1, 0) y–intercept: (0, 3) B. x–intercepts: (3, 0) and (1, 0) y–intercept: (0, 6) C. x–intercepts: (–3, 0) and (–1, 0) y–intercept: (0, 3) D. x–intercepts: (–3, 0) and (–1, 0) y–intercept: (0, 6)

OpenStudy (anonymous):

y = 2x2 – 8x + 6 = 2 (x-2)^2 -2 y = 2x2 – 8x + 6 = 2(x-3)(x-1) ans is B

OpenStudy (anonymous):

put y=0,you will get x-intercept x^2-4x+3=0 x^2-3x-x+3=0 x(x-3)-(x-3)=0 so x=1,x=3 simlarly y=6

OpenStudy (anonymous):

B